# FRISM: Fine-Grained Reasoning Injection via Subspace-Level Model Merging for Vision–Language Models

### v1

@@ 3.2 Theoretical Analysis
$\frac{\partial\mathcal{L}}{\partial\lambda_{i}}\approx h_{i}-2\alpha\|B_{i}\|_{F}^{2},$ (13)
@@
The detailed theoretical analysis is in the Appendix G.
@@ 4.1 Experimental Setup
Baseline Methods. We compare the proposed FRISM with baseline methods including: (1) Base VLM, i.e., the vision language model before merging; (2) Task Arithmetic (TA) (Ilharco et al., 2022); (3) Ties-Merging (Yadav et al., 2023); (4) DARE (Yu et al., 2024); and (5) IP-Merging (Hu et al., 2025). Please check Appendix C.2 for detailed information about baselines and hyper-parameter selection of existing merging methods.
Datasets. The performance of each model is evaluated by several multimodal tasks, including six vision-language reasoning tasks and three visual perception tasks. The vision-language reasoning tasks include: (1) MathVista (Lu et al., 2024a), (2) MathVision (Wang et al., 2024a), (3) MathVerse (Zhang et al., 2024), (4) MMMU (Yue et al., 2023), (5) R1-OneVision (Yang et al., 2025), and (6) MMStar (Chen et al., 2024a). The visual perception tasks include: (1) TextVQA (Singh et al., 2019), (2) POPE (Li et al., 2023b), (3) SeedBench (Li et al., 2023a). Please check Appendix C.3 for more details about the datasets. The evaluation details are declared in Appendix C.4. All the model IDs can be found in Appendix C.1.
4.2 Experimental Results
4.2.1 Experiments on Qwen2.5-VL series models
@@ 4.2 Experimental Results
Table 4: Ablation studies of FRISM.
@@
Results. The quantitative results are presented in Tab. 1. As shown, FRISM achieves consistent and significant improvements over existing merging methods across all model scales (3B, 7B, and 32B), demonstrating exceptional scalability and robustness. On the 3B scale, FRISM surpasses the best baseline method and the base VLM by 1.8%, while maintaining the visual perception score. Notably, the performance improvement on MathVista reaches 4.5%. This advantage is further amplified at the 7B scale, where FRISM obtains the highest reasoning accuracy of 49.4, surpassing the base VLM by 2.0% and outperforms all the baseline methods. It should be noted that the VL reasoning performance of FRISM on the 7B scale is close to ThinkLite-VL-7B (Wang et al., 2025b), which has undergone high-cost post-training. On VL perception tasks, existing methods show obvious performance degradation while FRISM shows a slight improvement compared to the base VLM. On the 32B scale, FRISM shows a more significant improvement of 2.4%. Notably, on MathVision, our method achieves an improvement of 4.4%. On VL perception tasks, the performance drop of FRISM is still the least among all the methods.
Keyword Analysis. To further validate the reasoning and vision capabilities of different models, we analyze the reasoning-related and vision-related keyword tokens generated by the models merged through different methods when merging Qwen2.5-VL-32B-Instruct and QwQ-32B on the MathVision dataset. Because QwQ-32B is a model capable of self-reflection and multi-step logical reasoning, we verify the strength of the reflective capabilities injected into different models by counting the number of reasoning-related reflection tokens in Fig. 5. It can be seen that FRISM generates the most reflection tokens compared to the base model and baseline methods, demonstrating that the reasoning capabilities are most effectively injected compared to other methods. Meanwhile, the number of vision-related tokens generated by FRISM is also improved compared to the base VLM, validating that FRISM preserves visual capabilities during merging. Please refer to Appendix J for more information.
@@ 4.2.2 Experiments on more types of models
Results. The results are in Tab. 2. As shown, FRISM demonstrates consistent and significant improvements, further validating its versatility and robustness. When merging ThinkLite-VL-7B with DeepSeek-R1-distilled-Qwen-7B, the VLM undergoes post-training for visual alignment and reasoning capabilities compared to pre-trained models. Consequently, the parameter differences between the VLM and LRM are significant under this setting, resulting in the failure of existing methods. None of the existing methods obtain performance gains through merging while FRISM achieves performance gains on nearly all tasks and retains visual performance on VL perception tasks. When merging Qwen2.5-VL-7B-Instruct with OpenThinker3-7B, FRISM achieves both the highest reasoning score and VL perception performance. In this setting, IP-Merging suffers from performance collapse, under $T=0.2$ and $T=0.3$. This issue is due to the exceptionally large merging coefficients calculated by IP-Merging and is discussed in detail in Appendix E. When applied to the InternVL3-8B architecture, FRISM also achieves remarkable performance on both VL reasoning and perception tasks. We also conduct experiments on Llama-based models in Appendix F.
@@ 4.4 Ablation Studies
Ablation on SVD. To further demonstrate the effectiveness of merging in the subspace level, in Tab. 4, we conduct an ablation study on the SVD process. Specifically, we compare our proposed FRISM with a variant that does not include the SVD process, denoted as FRISM (w/o SVD) in Tab. 4, a degenerated variant of adaptive layer-wise merging. Specifically, instead of fine-grained subspace modulation, we assign a single learnable gating scalar $\mathbf{g}^{(l)}\in\mathbb{R}$ to the entire task vector of the LRM. For this variant, the loss function and the training process follow FRISM. This variant yields a marginal improvement over the base VLM on VL reasoning benchmarks. In contrast, our FRISM significantly outperforms the variant without SVD, empirically demonstrating the advantages of merging in the subspace level over layer-wise merging.
@@ Appendix A Algorithm Flow of FRISM
$\begin{split}\mathbf{U},\mathbf{S},\mathbf{V}&=\mathrm{SVD}\left(\Delta\theta_{\text{lrm}}\right);\\ \Delta\theta_{\text{lrm}}^{\text{rank}=n}&=\mathbf{U}[:,n]\mathbf{S}[n,n]\mathbf{V}[n,:];\\ \theta_{\text{merged}}\left(n\right)&=\theta_{\text{vlm}}+\Delta\theta_{\text{lrm}}^{\text{rank}=n}\cdot\frac{\mathrm{\|\Delta\theta_{\text{lrm}}\|^{2}}}{\|\Delta\theta_{\text{lrm}}^{\text{rank}=n}\|^{2}}\cdot\lambda.\end{split}$ (14)
@@ C.1 Model IDs
In Table 5, we present the Huggingface (Wolf et al., 2019) IDs of the models used in our experiments, including VLMs and LRMs.
Table 5: Model-ID of the LLMs for the examined tasks.
@@
  $\quad\mathcal{L}_{\text{distill}}\leftarrow\mathrm{KLD}(\hat{\mathbf{y}}_{\text{vlm}}\|\hat{\mathbf{y}}_{\text{vlrm}})$; $\mathcal{L}_{\text{inject}}\leftarrow\|\mathrm{Sigmoid}(\boldsymbol{\mathbf{g}})\odot\mathbf{S}\|^{2}$
 $\quad\mathcal{L}\leftarrow\mathcal{L}_{\text{distill}}-\alpha\cdot\mathcal{L}_{\text{inject}}$
@@ C.2 Baseline Methods
Task Arithmetic (Ilharco et al., 2022) defines task vectors as the difference between finetuned model weights and the pre-trained model weights. Suppose a model $\theta_{i}$ is finetuned from a pre-trained model $\theta_{pre}$, the task vector is $\tau_{i}=\theta_{i}-\theta_{pre}$. When merging $\theta_{1..K}$, the merged model is $\theta_{M}=\lambda\sum_{i=1}^{K}\tau_{i}+\theta_{pre}$, where $\lambda$ is the merging coefficient. In our experiments, unless specified, we report the best performance among the merging coefficients among $\lambda\in[0.1,0.2,0.3]$.
@@
IP-Merging (Hu et al., 2025) is a merging method proposed specifically for transferring the math and reasoning capabilities from language reasoning models to multi-modal large language models. Directly merging a strong Math LLM into an VLM fails due to misalignment in parameter spaces and specific reasoning-associated layers. IP-Merging identifies crucial reasoning parameters and projects them into the VLM’s subspace to align them, allowing the VLM to directly absorb math reasoning abilities from an off-the-shelf Math LLM without losing its original capabilities. By setting a threshold, IP-Merging selectively merges only the parameter subspaces that exhibit similarity higher than the threshold after projecting them for alignment. Specifically, the higher threshold $T$ usually leads to fewer merged layers. In our experiments, unless specified, we report the best performance among the similarity threshold selections among $T\in[0.2,0.3,0.4]$.
@@ C.3 Datasets
R1-OneVision (Yang et al., 2025) is a large multimodal reasoning dataset with detailed step-by-step annotations across diverse domains, including mathematics, physics, chemistry, biology, and logical deduction, designed to train and evaluate models on complex cross-modal understanding and reasoning tasks. In our experiments, we apply the full dataset for evaluation.
@@ C.4 Evaluation
For VL reasoning datasets, we apply a step-by-step prompt. Box K shows a case of our prompt. For VL perception datasets, we use the default prompts provided by VLMEvalKit (Duan et al., 2024). During evaluation, all the models follow the greedy decoding strategy. We apply rule-based evaluation metrics for all the tasks.
Appendix D More results comparing subspace-level and layer-wise merging
Following the configuration of Fig. 3, we further plot the results by merging the VLMs and the task vectors from another model, OpenThinker3-7B (Guha et al., 2025), in Fig. 6. The results further demonstrate the effectiveness of merging in the subspace level.
Figure 6: Impact of scaling coefficients across subspace ranks when merging Qwen2.5-VL-7B-Instruct and the subspaces of OpenThinker task vector.
Appendix E Analysis of IP-Merging’s Failure Cases
$\lambda_{n}=\frac{\sum_{i=1}^{d}\sigma_{\text{vlm},i}^{n}}{\sum_{i=1}^{d}\sigma_{\text{lrm},i}^{n}},$ (15)
Table 6: Experiments of merging LLAVA-Next-8B with DeepThought-8B model.

### v2

@@ 3.2 Theoretical Analysis
$\frac{\partial\mathcal{L}}{\partial\lambda_{i}}\approx\left(h_{i}-2\alpha\|B_{i}\|_{F}^{2}\right)\lambda_{i},$ (13)
@@
A detailed theoretical analysis is in the Appendix F.
@@ 4.1 Experimental Setup
Baseline Methods. We compare the proposed FRISM with several baseline methods, including: (1) Base VLM, i.e., the vision language model before merging; (2) Task Arithmetic (TA) (Ilharco et al., 2023); (3) Ties-Merging (Yadav et al., 2023); (4) DARE (Yu et al., 2024); and (5) IP-Merging (Hu et al., 2025).
Please check Appendix C.2 for detailed information about baselines and hyper-parameter selection of existing merging methods.
@@
Datasets. The performance of each model is evaluated by several multimodal tasks, including six vision-language reasoning tasks and three visual perception tasks. The vision-language reasoning tasks include: (1) MathVista (Lu et al., 2024a), (2) MathVision (Wang et al., 2024a), (3) MathVerse (Zhang et al., 2024), (4) MMMU (Yue et al., 2023), (5) R1-OneVision (Yang et al., 2025b), and (6) MMStar (Chen et al., 2024a). The visual perception tasks include: (1) TextVQA (Singh et al., 2019), (2) POPE (Li et al., 2023b), (3) SeedBench (Li et al., 2023a). Please check Appendix C.3 for more details about the datasets. The evaluation details are declared in Appendix C.4. All the model IDs can be found in Appendix C.1.
4.2 Experimental Results
4.2.1 Experiments on Qwen2.5-VL series models
@@ 4.2 Experimental Results
Table 4: Ablation on the decomposition process and the injection loss of FRISM.
@@
Table 5: Ablation on the gate range of FRISM.
Methods MVista MMStar R1-OV TextVQA OCRBench
Qwen2.5-VL-7B-Inst 68.1 60.1 34.6 85.5 87.7
FRISM (Softplus) 68.8 60.3 36.0 84.2 87.4
FRISM (Sigmoid, Ours) 70.3 62.4 37.5 85.0 87.9
Table 6: Ablation on the decomposition strategy of FRISM.
Methods MVista MMStar R1-OV TextVQA OCRBench
Qwen2.5-VL-3B-Inst 40.5 40.5 28.0 79.3 82.6
FRISM (Rand-Mask) 41.3 41.9 27.8 79.0 82.3
FRISM (SVD, Ours) 45.0 43.4 28.9 79.2 82.7
Results. The quantitative results are presented in Tab. 1. As shown, FRISM achieves consistent and significant improvements over existing merging methods across all model scales (3B, 7B, and 32B), demonstrating exceptional scalability and robustness. On the 3B scale, FRISM surpasses the best baseline method and the base VLM by 1.8 points, while maintaining the visual perception score. Notably, the performance improvement on MathVista reaches 4.5 points. This advantage is further amplified at the 7B scale, where FRISM obtains the highest reasoning accuracy of 49.4, surpassing the base VLM by 2.0 points and outperforms all the baseline methods. It should be noted that the VL reasoning performance of FRISM on the 7B scale is close to ThinkLite-VL-7B (Wang et al., 2025b), which has undergone high-cost post-training. On VL perception tasks, existing methods show obvious performance degradation while FRISM shows a slight improvement compared to the base VLM. On the 32B scale, FRISM shows a more significant improvement of 2.4 points. Notably, on MathVision, our method achieves an improvement of 4.4 points. On VL perception tasks, the performance drop of FRISM is still the least among all the methods.
Keyword Analysis. To further validate the reasoning and vision capabilities of different models, we analyze the reasoning-related and vision-related keyword tokens generated by the models merged through different methods when merging Qwen2.5-VL-32B-Instruct and QwQ-32B on the MathVision dataset. Because QwQ-32B is a model capable of self-reflection and multi-step logical reasoning, we verify the strength of the reflective capabilities injected into different models by counting the number of reasoning-related reflection tokens in Fig. 5. It can be seen that FRISM generates the most reflection tokens compared to the base model and baseline methods, demonstrating that the reasoning capabilities are injected most effectively compared to other methods. Meanwhile, the number of vision-related tokens generated by FRISM is also improved compared to the base VLM, validating that FRISM preserves visual capabilities during merging. Please refer to Appendix I for more information. Visualization results are in Appendix G.
@@ 4.2.2 Experiments on more types of models
Results.
The results are in Tab. 2. As shown, FRISM demonstrates consistent and significant improvements, further validating its versatility and robustness. When merging ThinkLite-VL-7B with DeepSeek-R1-distilled-Qwen-7B, the VLM undergoes post-training for visual alignment and reasoning capabilities compared to pre-trained models. Consequently, the parameter differences between the VLM and LRM are significant under this setting, resulting in the failure of existing methods. None of the existing methods obtain performance gains through merging, while FRISM achieves performance gains on nearly all tasks and retains visual performance on VL perception tasks. When merging Qwen2.5-VL-7B-Instruct with OpenThinker3-7B, FRISM achieves both the highest reasoning score and VL perception performance. In this setting, IP-Merging suffers from performance collapse under $T=0.2$ and $T=0.3$. This issue is due to the exceptionally large merging coefficients calculated by IP-Merging and is discussed in detail in Appendix E. When applied to the InternVL3-8B architecture, FRISM also achieves remarkable performance on both VL reasoning and perception tasks. We also conduct experiments on Llama-based models in Appendix D.2.
@@ 4.4 Ablation Studies
Ablation on the decomposition process. To further demonstrate the effectiveness of merging in the subspace level, in Tab. 4, we conduct an ablation study on the SVD process. Specifically, we compare our proposed FRISM with a variant that does not include the SVD process, denoted as FRISM (w/o SVD) in Tab. 4, a degenerated variant of adaptive layer-wise merging. Specifically, instead of fine-grained subspace modulation, we assign a single learnable gating scalar $\mathbf{g}^{(l)}\in\mathbb{R}$ to the entire task vector of the LRM. For this variant, the loss function and the training process follow FRISM. This variant yields a marginal improvement over the base VLM on VL reasoning benchmarks. In contrast, our FRISM significantly outperforms the variant without SVD, empirically demonstrating the advantages of merging in the subspace level over layer-wise merging.
@@
Ablation on the gate range. We design the per-subspace Sigmoid gate as a bounded filter because post-hoc reasoning injection is fragile: unrestricted coefficients may cause visual drift, destabilize the merged model, and allow some subspaces to dominate, disrupting the model’s original properties. The overall injection magnitude is controlled by $\lambda_{\text{lrm}}$ and the gated singular values, while the sigmoid gate can provide conservative, fine-grained selection of reasoning subspaces under the visual-preservation constraint. An ablation study on the gate range is shown in Tab. 5. It can be seen that FRISM with the Sigmoid gate shows better performance compared to the variant with the Softplus gate.
Ablation on decomposition strategy. To validate the effectiveness of SVD compared to other decomposition methods, we apply random masks to LRM task vectors for decomposition. The results based on Qwen2.5-3B series models are shown in Tab. 6. It can be seen that FRISM (w/ SVD) consistently outperforms the variant with a random mask, further demonstrating the effectiveness of SVD. Additionally, compared to mask-based decomposition, subspaces are low-rank and require less GPU memory during training.
4.5 More Experimental Results
In Appendix D, we show additional experimental results, including results on more visual perception benchmarks and comparisons with more reasoning VLMs. We also present case studies in Appendix J and K.
@@ 5 Conclusions
Acknowledgements
This work is supported by National Key R&D Program of China (No. 2026YFE0101200), and Shanghai Natural Science Foundation (No. 23ZR1402900). The computations in this research were performed using the CFFF platform of Fudan University.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Algorithm Flow of FRISM
$\begin{split}\mathbf{U},\mathbf{S},\mathbf{V}&=\mathrm{SVD}\left(\Delta\theta_{\text{lrm}}\right);\\ \Delta\theta_{\text{lrm}}^{\text{rank}=n}&=\mathbf{U}[:,n]\mathbf{S}[n,n]\mathbf{V}[n,:];\\ \theta_{\text{merged}}\left(n\right)&=\theta_{\text{vlm}}+\Delta\theta_{\text{lrm}}^{\text{rank}=n}\cdot\frac{\mathrm{\|\Delta\theta_{\text{lrm}}\|^{2}}}{\|\Delta\theta_{\text{lrm}}^{\text{rank}=n}\|^{2}}\cdot\lambda.\end{split}$ (A1)
@@ C.1 Model IDs
In Table A1, we present the Huggingface (Wolf et al., 2019) IDs of the models used in our experiments, including VLMs and LRMs.
Table A1: Model-ID of the models, including VLMs and LRMs, for the examined tasks.
@@
DRIFT-VL-7B ChaoHuangCS/DRIFT-VL-7B
OpenVLThinker-7B ydeng9/OpenVLThinker-7B
VLAA-Thinker-Qwen2.5VL-7B UCSC-VLAA/VLAA-Thinker-Qwen2.5VL-7B
@@
  $\quad\mathcal{L}_{\text{distill}}\leftarrow\mathrm{KLD}(\hat{\mathbf{y}}_{\text{vlm}}\|\hat{\mathbf{y}}_{\text{vlrm}})$
 $\quad\mathcal{R}_{\text{inject}}\leftarrow\|\mathrm{Sigmoid}(\boldsymbol{\mathbf{g}})\odot\mathbf{S}\|^{2}$
 $\quad\mathcal{L}\leftarrow\mathcal{L}_{\text{distill}}-\alpha\cdot\mathcal{R}_{\text{inject}}$
 $\triangleright$ Let $\boldsymbol{\phi}=\{g^{\left(l\right)}\}^{L}_{l=1}$ denote all the learnable gates
@@ C.2 Baseline Methods
Task Arithmetic (Ilharco et al., 2023) defines task vectors as the difference between finetuned model weights and the pre-trained model weights. Suppose a model $\theta_{i}$ is finetuned from a pre-trained model $\theta_{pre}$, the task vector is $\tau_{i}=\theta_{i}-\theta_{pre}$. When merging $\theta_{1..K}$, the merged model is $\theta_{M}=\lambda\sum_{i=1}^{K}\tau_{i}+\theta_{pre}$, where $\lambda$ is the merging coefficient. In our experiments, unless specified, we report the best performance among the merging coefficients among $\lambda\in\{0.1,0.15,0.2\}$.
@@
IP-Merging (Hu et al., 2025) is a merging method proposed specifically for transferring the math and reasoning capabilities from language reasoning models to multi-modal large language models. Directly merging a strong Math LLM into an VLM fails due to misalignment in parameter spaces and specific reasoning-associated layers. IP-Merging identifies crucial reasoning parameters and projects them into the VLM’s subspace to align them, allowing the VLM to directly absorb math reasoning abilities from an off-the-shelf Math LLM without losing its original capabilities. By setting a threshold, IP-Merging selectively merges only the parameter subspaces that exhibit similarity higher than the threshold after projecting them for alignment. Specifically, the higher threshold $T$ usually leads to fewer merged layers. In our experiments, unless specified, we report the best performance among the similarity threshold selections among $T\in\{0.2,0.3,0.4\}$.
@@ C.3 Datasets
R1-OneVision (Yang et al., 2025b) is a large multimodal reasoning dataset with detailed step-by-step annotations across diverse domains, including mathematics, physics, chemistry, biology, and logical deduction, designed to train and evaluate models on complex cross-modal understanding and reasoning tasks. In our experiments, we apply the full dataset for evaluation.
@@ C.4 Evaluation
For VL reasoning datasets, we apply a step-by-step prompt. Box J shows a case of our prompt. For VL perception datasets, we use the default prompts provided by VLMEvalKit (Duan et al., 2024). During evaluation, all the models follow the greedy decoding strategy. We apply rule-based evaluation metrics for all the tasks.
Appendix D More Experimental Results
D.1 More experiments comparing with more models on more benchmarks
We further evaluate the reasoning and perception capabilities of the proposed FRISM. In Tab. A2, based on Qwen2.5VL-7B-Instruct model, we present additional comparisons with post-training methods, including an efficient post-training method, DRIFT-VL-7B (Huang et al., 2025), and two powerful post-trained models, OpenVLThinker-7B (Deng et al., 2025) and VLAA-Thinker-Qwen2.5VL-7B (Chen et al., 2025a). Additionally, we add experiments on three more visual perception benchmarks:
  • OCRBench (Liu et al., 2024b), evaluating OCR-related capabilities of multimodal large language models.
  • MME-Perception (Fu et al., 2026), evaluating perception-level abilities through a diverse set of visual understanding.
  • ChartQA (Masry et al., 2022), evaluating model’s ability to understand chart content.
The results in Tab. A2 show that FRISM is competitive with, and in our setting even slightly outperforms, several stronger post-training methods in average reasoning performance while preserving visual perception substantially better.
In Tab. A3, we compare the performance-cost trade-off of different methods. It can be seen that FRISM achieves impressive performance while requiring significantly lower computational resources. Additionally, FRISM only requires samples from unlabeled visual perception datasets for calibration, while other post-training methods usually require labeled VL reasoning datasets for training.
Table A2: Comparison of reasoning and vision performance across more methods on more benchmarks. The base model is Qwen2.5-VL-7B-Instruct. Note that Avg-R and Avg-V are respectively the reasoning and vision scores computed by first normalizing each benchmark score by the corresponding base VLM score and then averaging across benchmarks.
MVista MVerse MVision MMMU R1-OV MMStar Avg-R TextVQA POPE SeedBench MME OCRBench ChartQA Avg-V
Base 68.1 41.2 25.6 55.1 34.6 60.1 100% 85.5 86.4 77.0 1697.9 87.7 86.2 100%
Post-Training Methods
DRIFT 70.4 41.8 25.7 55.7 34.9 59.7 101.3% 84.1 84.3 70.7 1510.7 87.6 80.2 94.9%
OpenVLThinker-7B 69.4 41.5 26.5 54.1 35.0 62.5 101.7% 84.2 82.6 74.6 1411.6 83.4 84.2 94.5%
VLAA-Thinker 71.3 42.8 26.8 57.8 35.4 61.3 103.8% 82.9 86.2 75.2 1700.3 86.9 84.3 98.6%
Merging Methods
IP-Merging 67.2 40.8 25.5 55.9 36.0 60.7 100.6% 82.9 87.1 76.9 1694.8 88.4 81.6 98.8%
TaskArithmetic 65.4 37.8 24.2 53.7 36.6 60.7 97.9% 83.9 86.8 76.5 1674.9 88.5 72.9 97.0%
FRISM (Ours) 70.3 41.8 26.4 57.8 37.5 62.4 104.2% 85.0 87.1 76.9 1678.9 87.9 84.6 99.6%
Table A3: Comparison of training efficiency and performance.
Method #GPUs Training Time Avg-Reason Avg-Percep
Base 100% 100%
IP-Merging 100.6% 98.8%
DRIFT 8 $\sim$2 hours 101.3% 94.9%
OpenVLThinker-7B 8 >>1 day 101.7% 94.5%
VLAA-Thinker-7B 8 $\sim$6 hours 103.8% 98.6%
FRISM (Ours) 1 30 mins 104.2% 99.6%
Table A4: Experiments of merging LLAVA-Next-8B with DeepThought-8B model.